# Brain-inspired sparse training enables Transformers and LLMs to perform as fully connected

### v1

@@ 3.2 Sparse Topological Initialization with Bipartite Small-World model
Figure 3: The adjacency matrices of the Bipartite Scale-Free (BSF) network model and the Bipartite Small-World (BSW) network model vary with different values of $\beta$. a) The BSF model inherently forms a scale-free network characterized by a power-law distribution with $\gamma=2.76$. b) As $\beta$ changes from 0 to 1, the network exhibits reduced clustering. It is important to note that when $\beta=0$, the BSW model does not qualify as a small-world network.
We introduce all the network science based sparse topological initialization methods in Section C. As demonstrated in CHT (Zhang et al., 2024b), the final trained model converges into an ultra-small-world network, exhibiting both scale-free and small-world properties. This raises a compelling question: what if the network is directly initialized with these characteristics? In this work, we investigate this by initializing networks with bipartite small-world (BSW) and bipartite scale-free (BSF) structures. To further enhance the performance of BSF initialization, we introduce two optimization techniques. However, as detailed in Appendix D, our findings indicate that networks initialized with the BSW structure consistently achieve superior performance.
The BSW model, with its small-world properties, ensures both high clustering and a short average path length. The high clustering, which is absent in the BSF model, increases the probability that seed nodes share common neighbors along L3 paths (paths of length three). This, in turn, improves the effectiveness of the CH-based link predictor in generating accurate predictions early in the training process. In contrast, the BSF model’s preferential attachment mechanism creates hub nodes and a sparsely clustered structure at initialization, which is randomly assigned. Early in training, this random hub assignment may not align with the actual training dynamics, while the highly power-law degree distribution can mislead the CH link predictor, causing it to stack within the ELM.
3.3 Sigmoid Gradual Decrease Density
As demonstrated in GraNet (Liu et al., 2021) and MESTEM&S (Yuan et al., 2021), incorporating a density decrease strategy can significantly improve the performance of dynamic sparse training. In MESTEM&S, the density is reduced discretely at predefined epochs. In GraNet, the network evolution process consists of three steps: pruning, link removal, and link regrowth. The method first prunes the network to reduce the density, followed by removing and regrowing an equivalent number of links under the updated density. The density decrease in GraNet follows the same approach as Gradual Magnitude Pruning (GMP) (Zhu & Gupta, 2017), which adheres to a cubic function:
@@ 3.3 Sigmoid Gradual Decrease Density
$s_{t}=s_{f}+(s^{\prime}_{i}-s_{f})\left(\frac{1}{1+e^{-k\left(t-\frac{t_{f}-t_{0}}{2}\right)}}\right),$ (7)
where $k$ controls the curvature of the decrease. We set $k$=6 for all the experiments in this article. This strategy ensures a smoother initial pruning phase, allowing the model to warm up and stabilize before undergoing significant pruning, thereby enhancing training stability and performance.
Since our work focuses on MLP, Transformer, and LLMs, where FLOPs are linearly related to the density of the linear layers, the FLOPs of the whole training process are linearly related to the integral of the density function across the training time. the The integral of the GraNet decrease function from $t_{0}$ to $t_{f}$ is:
@@
In addition to refining the decay function, we replace the weight magnitude criterion used in the original GMP and GraNet processes with relative importance (RI). This adjustment is motivated by prior work (Zhang et al., 2024a), which has shown that RI provides a significant performance advantage over weight magnitude, particularly when pruning models initialized with high density.
4 Experiments
4.1 Experimental Setup
We evaluate the performance of CHTs using MLPs for image classification tasks on the MNIST (LeCun et al., 1998), Fashion MNIST (Xiao et al., 2017), and EMNIST (Cohen et al., 2017) datasets. To further validate our approach, we apply the sigmoid gradual density decay strategy to Transformers for machine translation tasks on the Multi30k en-de (Elliott et al., 2016), IWSLT14 en-de (Cettolo et al., 2014), and WMT17 en-de (Bojar et al., 2017) datasets. Additionally, we conduct language modeling experiments using the OpenWebText dataset (Gokaslan & Cohen, 2019) and evaluate zero-shot performance on the GLUE (Wang et al., 2019) and SuperGLUE (Wang et al., 2020) benchmark with LLaMA-130M (Touvron et al., 2023a). For MLP training, we sparsify all layers except the final layer, as ultra-sparsity in the output layer may lead to disconnected neurons, and the connections in the final layer are relatively minor compared to the previous layers. For Transformers and LLaMA-130M, we apply dynamic sparse training (DST) to all linear layers, excluding the embedding and final generator layer. Detailed hyperparameter settings for each experiment are provided in Tables 56, and 7.
Table 1: Performance comparison of different fixed sparsity dynamic sparse training methods on MNIST, Fashion MNIST (FMNIST), and EMNIST datasets trained on MLP at 99% sparsity. ACC represents accuracy, and ANP denotes the active neuron percolation rate that indicates the final size of the network. The best method for each dataset is highlighted in bold and the performances better than the fully connected ones are marked with “*”
Method MNIST FMNIST EMNIST
ACC (%) ANP (%) ACC (%) ANP (%) ACC (%) ANP (%)
FC 98.78 $\pm$ 0.02 - 90.88 $\pm$ 0.02 - 87.13 $\pm$ 0.04 -
CHTs (CH3-L3p) 98.81 $\pm$ 0.04* 20% 90.93 $\pm$ 0.03* 89% 87.61 $\pm$ 0.07* 24%
CHTs (CH2-L3n) 98.76 $\pm$ 0.05 27% 90.67 $\pm$ 0.05 73% 87.82 $\pm$ 0.04* 28%
CHT 98.48 $\pm$ 0.04 29% 88.70 $\pm$ 0.07 30% 86.35 $\pm$ 0.08 21%
RigL 98.61 $\pm$ 0.01 29% 89.91 $\pm$ 0.07 55% 86.94 $\pm$ 0.08 28%
SET 98.14 $\pm$ 0.02 100% 89.00 $\pm$ 0.09 100% 86.31 $\pm$ 0.08 100%
Baseline Methods
We compare our method with the baseline approaches in the literature. We divide the dynamic sparse training (DST) methods into two categories: fixed sparsity DST and density decay DST. For the fixed sparsiry DST, we compare CHTs with the SET (Mocanu et al., 2018), RigL (Evci et al., 2020), and CHT (Zhang et al., 2024b) and for the density decay DST methods, we compare CHTss with MESTEM&S (Yuan et al., 2021), GMP (Zhu & Gupta, 2017), and GraNet (Liu et al., 2021). We explain the detailed implementations and the reason we split GMP as a type of DST method in Appendix F.
4.2 MLP for image classification
Ablation Test.
Using MLP, we conduct an ablation study on each component proposed within the CHTs framework to determine the most effective implementation to apply next for the Transformer model. Figure 4a) compares the topologies initialized with the Bipartite Small-World (BSW) model at different values of $\beta$, clearly indicating that $\beta=0.25$ yields the best results. Figures 4b) assess the link removal methods, concluding that the weight magnitude soft (WM-soft) method outperforms all others. We consider the best settings showcased in these results to decide the CHTs strategy for training Transformers and LLaMA-130M.
Table 2: Performance comparison on machine translation tasks of Multi30k, IWSLT, and WMT with varying sparsity levels. The scores indicate BLEU scores, which is the higher the better. CHTs (GMP) indicates CHTs uses GMP’s density decay strategy. The best performance on each dataset is highlighted in bold and the performances better than the fully connected ones are marked with “*”. $s_{i}$ indicates the starting sparsity of the dst methods that use density decay strategy.
Method Multi30k IWSLT WMT
95% 90% 95% 90% 95% 90%
FC 31.28 24.2 25.22
SET 27.89 28.72 18.48 19.54 20.21 21.61
RigL 27.63 28.89 20.29 21.03 20.52 22.16
CHTs 28.12 30.35 20.55 21.60 21.14 22.68
MESTEM&S 28.71 28.26 18.95 20.77 20.79 22.3
GMP ($s_{i}=0.5$) 26.42 27.06 22.44 22.62 22.29 23.52
GraNet ($s_{i}=0.5$) 30.90 31.06 23.05 22.88 22.11 23.49
CHTs (GMP) ($s_{i}=0.5$) 30.49 30.33 23.68 23.64 22.8 23.22
CHTss ($s_{i}=0.5$) 32.04* 32.79* 24.86* 24.57* 22.68 24.05
Main Results.
In the MLP evaluation, we aim to assess the fundamental capacity of DST methods to train the fully connected module, which is common across many ANNs. The sparse topological initialization of CHTs is CSTI since the input bipartite layer can directly receive information from the input pixels. Table 1 displays the performance of DST methods compared to their fully connected counterparts across three basic datasets. The DST methods are tested at 99% sparsity. As shown in Table 1, both of the two regrowth methods of CHTs outperform the other fixed sparsity DST methods. Notably, the path-based CH3-L3p outperforms the fully connected one in all the datasets. The node-based CH2-L3n also achieves comparable performance on these basic datasets. In addition, we present the active neuron post-percolation rate (ANP) for each method in Table 1. It is evident that CHTs adaptively percolates the network more effectively while retaining performance.
4.3 Transformer on Machine Translation
We assess the Transformer’s performance on a classic machine translation task across three datasets. We take the best performance of the model on the validation set and report the BLEU on the test set. Beam search, with a beam size of 2, is employed to optimize the evaluation process. In our evaluation, CHTs configures the topology of each layer using the BSW model with $\beta=0.25$, employs WM-soft for link removal, and regrows new links using CH2-L3n-soft. Additionally, we apply an adjusted network percolation technique to the Transformer, as detailed in Appendix Section E. The findings, presented in Table 2, demonstrate that 1) CHTs surpasses other fixed density DST methods on all the sparsity scenrios. 2) Incorperating with the sigmoid density decrease, CHTss outperforms even the fully connected ones with only 5% density.
Table 3: Validation perplexity of different dynamic sparse training (DST) methods on OpenWebText using LLaMA-130M across varying sparsity levels. Bold values denote the best performance among DST methods. Lower perplexity corresponds to better model performance. $s_{i}$ represents the initial sparsity for DST methods employing a density decay strategy.
Method Sparsity
95% 90% 80% 70%
FC 19.27
SET 28.37 24.73 22.02 20.82
RigL 49.39 37.18 66.35 25.85
CHTs 27.72 24.24 21.70 21.15
${}_{\text{EM\&S}}$ 27.96 24.98 22.21 21.32
GMP ($s_{i}=0.5$) 27.16 23.61 22.28 20.49
GraNet ($s_{i}=0.5$) 61.31 26.81 29.03 22.84
CHTs (GMP) ($s_{i}=0.5$) 26.81 22.94 20.94 20.01
CHTss ($s_{i}=0.5$) 25.29 22.71 20.78 19.92
4.4 Natural Language Generation

### v2

@@ Appendix D Sparse topological initialization
Figure 5: The adjacency matrix of the Bipartite Scale-Free (BSF) network model compared to those of the Bipartite Small-World (BSW) network, the Bipartite Receptive Field with fixed sampling (BRFf), and the Bipartite Receptive field with uniform sampling (BRFu) as parameters $\beta$ and $r$ vary between 0 and 1. a) The BSF model inherently forms a scale-free network characterized by a power-law distribution with $\gamma=2.76$. b) As $\beta$ changes from 0 to 1, the network exhibits reduced clustering. It is important to note that when $\beta=0$, the BSW model does not qualify as a small-world network. c) As $r$ increases towards $1$, the adjacency matrix becomes more random, while sampling the output neurons’ degrees from a fixed or uniform distribution.
Bipartite receptive field model.
The Bipartite Receptive Field (BRF) model is a random network generation technique designed to mimic the receptive field phenomenon in the brain networks. The process involves adding links to the adjacency matrix of the bipartite network, with the connectivity structured around the main diagonal according to a parameter $r\in[0,1]$. A low value of $r$ results in links that are primarily clustered around the diagonal, while a higher value of $r$ leads to a more random connectivity pattern. Specifically, a bipartite adjacency matrix with links near the diagonal indicates that adjacent nodes from the two layers are linked, whereas links far from the diagonal correspond to more distant node pairs.  Mathematically, consider an $N\times M$ bipartite adjacency matrix ${M_{i,j}}_{i=1,\dots,M,j=1,\dots,N}$, where $M$ represents the input size and $N$ represents the output size. Each entry of the matrix $m_{i,j}$ is set to $1$ if input node $i$ is connected to output node $j$, and 00 otherwise. A scoring function $S_{i,j}$ is assigned to each connection in the adjacency matrix based on its distance to the main diagonal. This score is given by:
$S_{i,j}=d_{ij}^{\frac{1-r}{r}},$ (7)
where
$d_{ij}=min\{|i-j|,|(i-M)-j|,|i-(j-N)|\}$ (8)
is the distance between the input and output neurons. Therefore, $S_{i,j}$ represents the distance from the diagonal, raised to the power of $\frac{1-r}{r}$. The parameter $r$ controls how structured or random the adjacency matrix is. As $r\to 0$, the scoring function becomes more deterministic, with high scores assigned to entries near the diagonal and low scores to entries farther away. Conversely, as $r\to 1$, all scores $S_{i,j}$ become more uniform, leading to a more random, less structured adjacency matrix. The next step is to determine the degree distribution for the output nodes. This can either be fixed, assigning the same degree to all output nodes, or uniform, where the degrees are randomly sampled from a uniform distribution. Hence, we propose two variations of the BRF model: the Bipartite Receptive Field with fixed sampling (BRFf), in which the degrees of output nodes are fixed, and the Bipartite Receptive Field with uniform sampling (BRFu), where the degrees of the output nodes follow a uniform distribution. This represents an additional enhancement to the WS scheme, which offers no way to control how connections are allocated among the output nodes. In conclusion, to run the BRF model, the user should input an output degree distribution and a spatial dependent distance randomness.
Figure 6: The ablation test of the $\beta$ of the bipartite small world (BSW) model and the removal methods in CHTs. a) evaluates the influence of the rewiring rate $\beta$ on the model performance when initialized with the BSW model. b) assesses the influence of link removal selecting from the weight magnitude (WM), weight magnitude soft (WM-soft), relative importance (RI), and relative importance soft (RI-soft). We utilize the win rate of the compared factors under the same setting across each realization of 3 seeds for all experiment combinations on MLP. The factor with the highest win rate is highlighted in orange.
Figure 7: The ablation test of the parameter $r$ in the bipartite receptive field (BRF) model and the removal methods in CHTs using the BRF initialization technique. a) evaluates the influence of the parameter $r$ on the model performance when initialized with the BRF model. b) assesses the influence of link removal in the CHTs model with BRF initialization. We utilize the win rate of the compared factors under the same setting across each realization of 3 seeds for all experiment combinations on MLP. The factor with the highest win rate is highlighted in orange.
Appendix E Equal Partition and Neuron Resorting to enhance bipartite scale-free network initialization
As indicated in SET and CHT [5, 9], trained sparse models typically converge to a scale-free network. This suggests that initiating the network with a scale-free structure might initially enhance performance. However, starting directly with a Bipartite Scale-Free model (BSF, power-law exponent $\gamma=2.76$) does not yield effective results. Upon deeper examination, two potential reasons emerge:
  • As demonstrated in CHT, in the final network, the hub nodes of one layer’s output should correspond to the input layer of the subsequent layer, which means the hub nodes should have a high degree on both sides of the layer. However, the BSF model’s random selection disrupts this correspondence, significantly reducing the number of Credit Assignment Paths (CAP) [9] in the model. CAP is defined as the chain of transformation from input to output, which counts the number of links that go through the hub nodes in the middle layers.
To address these issues, we propose two solutions:
Figure 8: The Performance of the bipartite scale-free model and two enhanced techniques. a) shows the win rate of the Bipartite Scale-Free network model (BSF) with the different techniques. EP stands for equal partition of the first layer, and Resort refers to reordering the neurons based on their degree. b) assesses the comparison between Correlated Sparse Topological Initialization (CSTI), the Bipartite Scale-Free (BSF) model with the best solution from a), and the Bipartite Small-World (BSW) model with $\beta=0.25$.
Appendix F Network percolation and extension to Transformer.
We have adapted network percolation [53, 9] to suit the architecture of the Transformer after link removal. The core idea is to identify inactive neurons, which are characterized by having no connections on either one or both sides within a layer of neurons. Such neurons disrupt the flow of information during forward propagation or backpropagation. In addition, Layer-wise computation of the CH link prediction score further implies that neurons without connections on one side are unlikely to form connections in the future. Therefore, network percolation becomes essential to optimize the use of remaining links.
As shown in Figure 1, network percolation encompasses two primary processes: c1) inactive neuron removal to remove the neurons that lack connections on one or both sides; c2) incomplete path adjustment to remove the incomplete paths where links connect to the inactive neurons after c1). Typically applied in simpler continuous layers like those in an MLP, network percolation requires modification for more complex structures. For example, within the Transformer’s self-attention module, the outputs of the query and key layers undergo a dot product operation. It necessitates percolation in these layers to examine the activity of the neurons in both output layers at the same position. Similar interventions are necessary in the up_proj and gate_proj layers of the MLP module in the LLaMA model family [1, 54].
Appendix G Baseline Methods
G.1 Fixed Density Dynamic Sparse Training Methods
SET
[5]: Removes connections based on weight magnitude and randomly regrows new links.
RigL
[7]: Removes connections based on weight magnitude and regrows links using gradient information, gradually reducing the proportion of updated connections over time.
CHT
[9]: A state-of-the-art (SOTA) gradient-free method that removes links with weight magnitude and regrows links based on CH3-L3 scores. CHT is often applied with early stopping to mitigate its computational complexity when working with large models.
G.2 Gradual Density Decrease Dynamic Sparse Training Methods
GMP
[55, 43]: Prunes the network with weight magnitude and gradually decreases the density based on Equation 10. Although originally a pruning method, GMP is treated as a dynamic sparse training method in their implementation [43], as it stores historical weights and allows pruned weights to reappear during training, since, during training, the pruning threshold might change.
MESTEM&S
[8]: Implements a two-stage density decrease strategy as described in the original work. It removes links based on the combination of weight magnitude and 0.01*gradient and regrows new links randomly.
GraNet
[42]: Gradually decreases density using Equation 10. Similar to RigL, GraNet removes links based on the weight magnitude and regrows new links with the gradient of the existing links.
Table 11: Float32 Precision Comparison on LLaMA-130M. Bold values denote the best performance among DST methods. Lower perplexity corresponds to better model performance. $s_{i}$ represents the initial sparsity for DST methods employing a density decay strategy.
Method Sparsity
70% 80%
FC 17.07
RigL 18.34 19.64
CHTs 17.99 19.25
GraNet ($s_{i}=0.5$) 17.92 18.79
CHTss ($s_{i}=0.5$) 17.76 18.69
Appendix H Ablation and Sensitivity Tests
Table 12: Ablation results of Transformer on Multi30K and IWSLT datasets at 90% sparsity. The scores indicate BLEU scores, the higher the better. Bold values denote the best performance among DST methods.
Variant Multi30K (90% sparsity) IWSLT (90% sparsity)
a. CHTss 28.38 19.91
b. CHTss    without node-based implementation 32.68 (2.42 hours) 24.82 (18 hours)
c. CHTss    without soft sampling 28.92 21.88
d. CHTss    without sigmoid decay (= CHTs) 30.35 21.60
e. CHTss    (full model) 32.79 (0.25 hours) 24.57 (1.5 hours)
An overall ablation test
To fully assess each component’s effectiveness, we conduct several ablation and sensitivity tests that help us understand how to select a sparse topological initialization and identify the best link removal and regrowth methods. We first made a global test for all the components in Table 12, which shows the effectiveness of each element introduced by this article. The node-based and path-based link regrowth methods have comparable performance, but the node-based versions are much faster.
Sparse topological initialization.
For sparse topological initialization, we compare BRF, BSW, BSF, and CSTI [9] across three image classification datasets, as shown in Figure 8b. The results indicate that when the inputs can directly access task-relevant information, CSTI consistently achieves the best performance. In general, BRF and BSW perform similarly under these conditions, but outperform the BSF initialization.
To further validate our findings, we evaluate BRF and BSW network initializations on machine translation tasks using Transformer models. Figure 9 and Figure 10 present the performance comparisons between BSW and BRF on the Multi30k and IWSLT datasets, while Figure 11 shows the win-rate analysis. These comparisons demonstrate that BRF consistently outperforms BSW across most cases. Additionally, Figure 7a analyzes the impact of the receptive field range $r$ on BRF initialization for MNIST, Fashion MNIST, and EMNIST tasks using MLPs, with results indicating that $r=0.25$ yields the best performance.
Building on this prior knowledge, we further evaluate BRF on LLaMA-60M and LLaMA-130M models, testing $r$ values in the range $[0,0.3]$ and comparing two different degree distributions. The results, shown in Table 9 and Table 10, indicate that on LLaMA models, the choice of $r$ and distribution has limited impact. While $r=0.1$ wins slightly more often, the improvements remain marginal. Finally, Table 3 reports the best performance combinations of $r$ and degree distributions derived from these evaluations.
Table 13: Performance comparison of CHTs and CHTss at 90% sparsity across different removal methods. The tested dataset is Multi30K, and the reported metric is BLEU, which is the higher the better.
Remove Method CHTs CHTss
set 28.82 25.76
wm 28.17 31.15
wm_soft 30.35 32.79
ri 28.91 32.20
ri_soft 27.86 31.86
MEST 28.70 32.07
snip 28.23 31.66
sensitivity 29.02 29.73
Rsensitivity 28.18 30.67
Link removal.
We first conduct a simple evaluation of the link removal methods introduced in this article when changing the $\alpha$ and $\delta$ inside Figure 1b2) on Figure 6b) and Figure 7b). The removal methods are selected from Weight Magnitude (WM), Weight Magnitude soft (WMs), Relative Importance (RI), and Relative Importance soft (RIs). For WM we fix the hyperparameters $\alpha=1$ and $\delta=1$; for RI we fix the hyperparameters $\alpha=0$ and $\delta=0.5$; for WMs we fix $\alpha=1$ and we let $\delta$ increase linearly from $0.5$ to $0.9$; for RIs we fix $\alpha=0$ and let $\delta$ increase linearly from $0.5$ to $0.9$. From the results, it can be observed that WMs performs the best in most cases. We compare these methods with those in [56] in Table 13 on two machine translation tasks. The results indicate that using WMs as a link removal method generally outperforms the alternatives.
We also evaluate how to define the softness in WMs. During sampling, we have a hyperparameter to decide the temperature of the scores that convert to the probability of being removed. We perform a test using a linear decay solution, since, generally, the weights in the model become more reliable as training progresses. Figure 12 shows the variation in BLEU scores as we change the starting and ending values of the $\delta$ parameter in the soft weight magnitude removal method on transformer models. Recalling that we define the temperature by $T=\frac{1}{1-\delta}$, we observe that for a simple benchmark like Multi30k, a high starting temperature produces better performance. This is motivated by the fact that loss decreases very fast through epochs, meaning that weights are learned quickly, and we can deterministically remove weights with high reliability. In more complex datasets, like IWSLT, low starting temperatures are preferred. This is because during the early stages of training, weights are learned slowly, meaning that a deterministic removal can be less reliable. To be more consistent, we select a start $\delta=0.5$ and end $\delta=0.9$ for all the tasks in the main article.
Figure 9: Top average BLEU for WS (grey) and BRF initialization methods on the Multi30k translation dataset of CHTs (left) and CHTss (right, with sigmoid decay) at different sparsity levels. Error bars denote the standard error across three seeds.
Figure 10: Top average BLEU for WS (grey) and BRF initialization methods on the IWSLT translation dataset of CHTs (left) and CHTss (right, with sigmoid decay) at different sparsity levels. Error bars denote the standard error across two seeds.
Figure 11: Win rates of BRF against WS over CHTs and CHTss models on different datasets (Multi30k and IWSLT) and different sparsities (0.9 and 0.95 for IWSLT and 0.7, 0.8, 0.9, 0.95 for Multi30k).
Figure 12: Investigating the level of randomness in link removal strategies. Top BLEU scores of the transformer model using CHTs with weight magnitude soft removal strategy, as the initial and final values of $\delta$ take values in $\{0.1,0.25,0.5,0.75,0.8\}$.
Appendix I Integral of Sigmoid Density Decay and Cubic Density Decay
In this section, we show the formula for the proposed Sigmoid Density Decay and the Cubic decay implemented in GraNet [42] and GMP [55].
For the Cubic function, it is formulated as:
@@ Appendix I Integral of Sigmoid Density Decay and Cubic Density Decay
Since our work focuses on MLP, Transformer, and LLMs, where FLOPs are linearly related to the density of the linear layers, the FLOPs of the whole training process are linearly related to the integral of the density function across the training time. The integral of the cubic decay function from $t_{0}$ to $t_{f}$ is:
@@
Appendix J Historical weights
Inspired by GMP [55, 43], we incorporate historical weights into our CHTs and CHTss implementation. During training, we maintain a historical weight matrix that records previously learned weights throughout the training process. When CHTs and CHTss predict new links, we initialize them using their corresponding historical weights - specifically, the values they held before being pruned. In this way, CHTs and CHTss enable weight recovery with preserved memory, allowing the model to retain valuable prior information.
Appendix K Pseudo code of node-based CH link predictor
We present the pseudocode for the node-based CH2-L3 regrowth method in Algorithm 1, which comprises three main steps.
Step 1: We compute the sets of path length-2 (L2) neighbors from nodes $U$ to $U$ and from $V$ to $V$. The computational complexity of this step is
$\mathcal{O}(m\left\langle d_{m}d_{n}\right\rangle+n\left\langle d_{m}d_{n}\right\rangle),$
where $\left\langle d_{m}d_{n}\right\rangle$ denotes the average product of degrees $d_{m}$ and $d_{n}$ across all relevant node pairs. In the case of an ultra-sparse network (e.g., with average degree close to 1), this simplifies to $\mathcal{O}(m+n)$.
Step 2: We treat each destination node along the $UU$ and $VV$ paths as a potential common neighbor and compute the CH2-L3 score according to Equation (3) in the main text. This step has a time complexity of
$\mathcal{O}(m^{2}+n^{2}).$
Step 3: We aggregate the CH2-L3 scores across all hop-3 nodes. This step requires
$\mathcal{O}(mn\left\langle d_{m}\right\rangle+nm\left\langle d_{n}\right\rangle)$
time. Under the ultra-sparse assumption, this further reduces to $\mathcal{O}(mn+nm)$.
Combining all steps, the overall time complexity of the node-based CH2-L3 regrowth procedure is
$\mathcal{O}(mn\left\langle d_{m}\right\rangle+nm\left\langle d_{n}\right\rangle).$
Since all operations can be implemented in a matrix-wise manner leveraging GPU acceleration assuming all matrices as dense, the total time complexity becomes
$\mathcal{O}(nm^{2}+mn^{2}).$
Algorithm 1 CH2–L3n
1:  Binary bipartite adjacency matrix $A_{UV}\!\in\!\{0,1\}^{m\times n}$triangleright $U$–to–$V$ edges
2:  Score matrix $S\!\in\!\mathbb{R}^{m\times n}$ with $S[i,j]>0$ only if $A_{UV}[i,j]=0$
3:  function CH2-L3n($A_{UV}$)
4:       $A_{VU}\leftarrow A_{UV}^{\top}$triangleright $V\!\to\!U$ edges
5:       $d_{U}\leftarrow$ row-sum$(A_{UV})$
6:       $d_{V}\leftarrow$ row-sum$(A_{VU})$
7:  $\triangleright$ Step 1: two–step paths inside each partition
8:       $UU\leftarrow A_{UV}A_{VU}$triangleright $U\!\to\!V\!\to\!U$
9:       $VV\leftarrow A_{VU}A_{UV}$triangleright $V\!\to\!U\!\to\!V$
10:  $\triangleright$ Step 2: CH2-L3n preparatory scores
11:       init $e_{UU}\leftarrow 0_{m\times m},\;e_{VV}\leftarrow 0_{n\times n}$
12:       for $i\leftarrow 1$ to $m$ do
13:           for $j\leftarrow 1$ to $m$ s.t. $j\neq i$ and $UU[i,j]>0$ do
14:                $ext\leftarrow d_{U}[j]-UU[i,j]-1$
15:                $e_{UU}[i,j]\leftarrow\dfrac{UU[i,j]+1}{ext+1}$triangleright # According to Equation (3)
16:           end for
17:       end for
18:       for $a\leftarrow 1$ to $n$ do
19:           for $b\leftarrow 1$ to $n$ s.t. $b\neq a$ and $VV[a,b]>0$ do
20:                $ext\leftarrow d_{V}[b]-VV[a,b]-1$
21:                $e_{VV}[a,b]\leftarrow\dfrac{VV[a,b]+1}{ext+1}$triangleright # According to Equation (3)
22:           end for
23:       end for
24:  $\triangleright$ Step 3: final CH2–L3n scores
25:       init $S\leftarrow 0_{m\times n}$
26:       for $i\leftarrow 1$ to $m$ do
27:           for $a\leftarrow 1$ to $n$ s.t. $A_{UV}[i,a]=0$ do
28:                $S_{UV}\leftarrow\sum\limits_{j=1}^{m}e_{UU}[i,j]\;A_{UV}[j,a]$
29:                $S_{VU}\leftarrow\sum\limits_{b=1}^{n}e_{VV}[a,b]\;A_{VU}[b,i]$
30:                $S[i,a]\leftarrow S_{UV}+S_{VU}$
31:           end for
32:       end for
33:       return $S$
34:  end function
Appendix L Extra results of LLaMA1b